Kareem purchased a leather jacket for $800 with his new credit card. He budgeted $50 a month to pay toward the debt. When he read the fine print on the credit card statement, he saw that for each month he owed a balance, a finance charge of $17.50 would be added. How many months will it take Kareem to pay off the leather jacket?
Thanks
--------------------------------------------------------------------------------
One machine in a print shop can produce a certain # of pages twice as fast as another machine. Operating together these machines can produce this number of pages in 8 minutes. How long would it take the slower machine working alone to produce this number of pages?
______minutes.
Thanks

Answer by stanbon(48569)

Kareem purchased a leather jacket for $800 with his new credit card. He budgeted $50 a month to pay toward the debt. When he read the fine print on the credit card statement, he saw that for each month he owed a balance, a finance charge of $17.50 would be added. How many months will it take Kareem to pay off the leather jacket?
Let number of months needed to pay off the jacket be "x".
PAYING Data:
$50 per month for x months = 50x Dollars
OWING Data:
$800 + $17.5x
EQUATION:
50x=800+17.5x
32.5x=800
x=24.615....
Number of months will be 1 more than 24 or 25 months.
------------------------------------------------------
One machine in a print shop can produce a certain # of pages twice as fast as another machine. Operating together these machines can produce this number of pages in 8 minutes. How long would it take the slower machine working alone to produce this number of pages?
______minutes.

SLOWER MACHINE DATA:
Number of minutes for job= x; Rate of production = 1/x job/min
FASTER MACHINE DATA:
Number of minutes for job= 2x; Rate of production = 1/2x job/min
TOGETHER DATA:
Number of minutes fo job= 8; Rate of production = 1/8 job/min
EQUATION:
slower rate + raster rate = together rate
1/x + 1/2x = 1/8
Multiply through by 8x to get:
8+4=x
x=12 minutes (time for slower machine to do the job)
Cheers,
Stan H.